OpenStudy (shelby1290):

A rectangle has a width of 2x-3 and a length of 3x+1 a) Write its area as a simplified polynomial b) Write expressions for the dimensions if the width is doubled and the length is increased by 2 c) Write the new area as a simplified polynomial

OpenStudy (anonymous):

Where are you having issues with this problem?

OpenStudy (shelby1290):

@LifeEngineer I don't know what to do

OpenStudy (shadowlegendx):

What is the formula for the area of a triangle

OpenStudy (shadowlegendx):

*rectangle

OpenStudy (shelby1290):

base times height

OpenStudy (shadowlegendx):

Area = Length times Width

OpenStudy (shadowlegendx):

So multiply the length, by the width

OpenStudy (shelby1290):

Whoops! Yea, length times width is what I meant to say

OpenStudy (shelby1290):

(2x-3)(3x+1) like that?

OpenStudy (alexandervonhumboldt2):

yes now foil this

OpenStudy (shadowlegendx):

Length = 3x+1 Width = 2x-3 Area = (2x-3)(3x+1) FOIL First, Outer, Inner, Last First = \[2x \times 3x\] First = 6x Outer = \[2x \times 1 \] Outer = 2x Inner = \[-3 \times 3x\] Inner = -9x Last = -3 x 1 Last = -3

OpenStudy (shadowlegendx):

So now we add them all up First is actually 6x^2

OpenStudy (shadowlegendx):

\[6x^2 + 2x - 9x - 3\] What do we get?

OpenStudy (shelby1290):

(2x-3)(3x+1) =6x^2+2x-9-3 =6x^2+2x-12

OpenStudy (shadowlegendx):

thatʻs not right

OpenStudy (shadowlegendx):

Itʻs -9x , not -9

OpenStudy (shelby1290):

(2x-3)(3x+1) =6x^2+2x-9x-3 =6x^2-7x-3

OpenStudy (shadowlegendx):

Correct

OpenStudy (shadowlegendx):

We have solved a)

OpenStudy (shelby1290):

Is that the area?

OpenStudy (shadowlegendx):

Yes

OpenStudy (shelby1290):

Does that mean that I have to put cm^2 at the end of it?

OpenStudy (shadowlegendx):

If that is the unit of measure, then yes

OpenStudy (shadowlegendx):

Now b) Write expressions for the dimensions if the width is doubled and the length is increased by 2 Width = 2x-3 Length = 3x+1 Doubled is multiplication We basically multiply 2x-3 by itself This is represented by... (2x - 3)^2 which is basically (2x - 3)(2x - 3) So we foil again

OpenStudy (shadowlegendx):

Can you try do that one?

OpenStudy (shelby1290):

(2x-3)(2x-3) =4x^2-6x-6x+9 =4x^2-12x+9

OpenStudy (shadowlegendx):

Correct, now that is the doubled form of 2x - 3

OpenStudy (shadowlegendx):

Now we just increase 3x+1 by 2

OpenStudy (shadowlegendx):

Do you know what "increased by" means?

OpenStudy (shelby1290):

(3x+1)(3x+1) =9x^2+3x+3x+1 =9x^2+6x+1

OpenStudy (shadowlegendx):

No...that is doubling it

OpenStudy (shadowlegendx):

We donʻt want to multiply it by itself, we want to do a different operation with the number 2

OpenStudy (shelby1290):

What does increased by 2 mean? To the power of two?

OpenStudy (shelby1290):

(3x+1)^2 like that?

OpenStudy (shadowlegendx):

no

OpenStudy (shadowlegendx):

Let says I have a laser gun. It is currently at power level two. If I increase the power level by 2, what level will it be at?

OpenStudy (shelby1290):

4 ?

OpenStudy (shadowlegendx):

Yes, but what mathematical operation would you do to get there?

OpenStudy (shelby1290):

addition?

OpenStudy (shadowlegendx):

Yes

OpenStudy (shadowlegendx):

So 3x + 1 increased by 2

OpenStudy (shelby1290):

3x+1 turns into.... 5x+3

OpenStudy (shadowlegendx):

How did you get that

OpenStudy (shelby1290):

3x +2 = 5x 1 +2 = 3

OpenStudy (shadowlegendx):

No... You canʻt do 3x + 2 Do you know about like terms?

OpenStudy (shadowlegendx):

2 does not have the variable x, thus it is not like terms with 3x, and cannot be added or subtracted from each other

OpenStudy (shadowlegendx):

They can, however, multiply and divide each other

OpenStudy (shelby1290):

Yes, I'm familiar with like terms. I'm just confused.

OpenStudy (shadowlegendx):

So 3x + 1 increased by 2 What do we get? ^-^

OpenStudy (shelby1290):

3x times 2 = 6x 1 times 2 = 2

OpenStudy (shelby1290):

6x+2

OpenStudy (shadowlegendx):

noooo We are doing addition!

OpenStudy (shelby1290):

But you said: "2 does not have the variable x, thus it is not like terms with 3x, and cannot be added or subtracted from each other"

OpenStudy (shelby1290):

I thought we didn't have to do addition

OpenStudy (shadowlegendx):

Exactly BUT increased by = + So 3x + 1 + 2

OpenStudy (shadowlegendx):

We donʻt add the 2 to 3x.

OpenStudy (shelby1290):

So would 3x + 1 + 2 be my final answer or would it be 3x+3 ?

OpenStudy (shadowlegendx):

Yes

OpenStudy (shelby1290):

Okay and what's the new area?

OpenStudy (shadowlegendx):

c) Write the new area as a simplified polynomial Our doubled width = 4x^2-12x+9 Our new length = 3x + 3

OpenStudy (shadowlegendx):

So we have to multiply those two new dimensions

OpenStudy (shelby1290):

(4x^2-12x+9)(3x+3) Do I have to foil this?

OpenStudy (shadowlegendx):

We cannot FOIL this (4x^2-12x+9)(3x+3) First we start by pulling out like factors 3x + 3 = 3 • (x + 1) or 3(x + 1)

OpenStudy (shadowlegendx):

Next, we factor by splitting the middle term

OpenStudy (shadowlegendx):

Factoring 4x2 - 12x + 9 The first term is, 4x2 its coefficient is 4 . The middle term is, -12x its coefficient is -12 . The last term, "the constant", is +9 Multiply the coefficient of the first term by the constant 4 • 9 = 36 Find two factors of 36 whose sum equals the coefficient of the middle term, which is -12 . -36 + -1 = -37 -18 + -2 = -20 -12 + -3 = -15 -9 + -4 = -13 -6 + -6 = -12

OpenStudy (shadowlegendx):

Rewrite the polynomial splitting the middle term using the two factors found in the above post -6 4x2 - 6x - 6x - 9 Add up the first 2 terms, pulling out like factors 2x • (2x-3) Add up the last 2 terms, pulling out common factors 3 • (2x-3) Add up the four terms of previous step (2x-3) • (2x-3) Which is the desired factorization. We multiply (2x-3)(2x-3) We get = 3 • (2x - 3)2 • (x + 1)

OpenStudy (shadowlegendx):

Correction: 3(2x - 3)^2 • (x + 1)

OpenStudy (shadowlegendx):

Simplified that is 12x^3 -24x^2 - 9x + 27

OpenStudy (shadowlegendx):

That is your new area